OpenStudy (anonymous):

|Use factoring to solve this equation.| x^2+x=6 What is x?

jimthompson5910 (jim_thompson5910):

x^2+x=6 x^2+x-6 = 0 (x+3)(x-2) = 0 I'll let you finish

jimthompson5910 (jim_thompson5910):

To factor x^2+x-6, you find two numbers that multiply to -6 and add to 1 at the same time. These two numbers are 3 and -2, so that's why it factors to (x+3)(x-2)

OpenStudy (anonymous):

Thank you good sir~

jimthompson5910 (jim_thompson5910):

you're welcome